OpenStudy (anonymous):

Quick math question. Please help asap Apples cost $5 and orangs cost $3. In a bag full of apples and oranges, the total cost is $85. What is the number of oranges in the bag

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

so tired

OpenStudy (anonymous):

i know so far that it is 5x + 3y = 85 but thats it

OpenStudy (anonymous):

10 oranges = 30 + 11 apples = 85

OpenStudy (anonymous):

10 oranges = $30 and 11 apples = 55 dollars

OpenStudy (danjs):

how many total apples and oranges are there x + y =

OpenStudy (anonymous):

so that is $85

OpenStudy (anonymous):

can I get medal?

OpenStudy (anonymous):

there are 10 oranges

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

@DanJS thats why im confused. if i knew the total amount then i would be able to solve the problem through a equation

OpenStudy (anonymous):

@Liam2004 thnx Liam. Your right because i have the answer with me but i was hoping that there was a way to do it through an equation

OpenStudy (anonymous):

o = orange and a = apple

OpenStudy (anonymous):

so 10o + 11a = $85

OpenStudy (danjs):

what if you had 5 oranges instead and 14 apples

OpenStudy (danjs):

you have enough for 1 equation / 2 variables need another equation, i am guessing the total number of apples and oranges in the bag is somewhere so you can say x + y = (total in bag) then solve the two equations

OpenStudy (danjs):

It says, 'in a bag full', that seems like a specific amount of fruit , idk, is this multi part

OpenStudy (anonymous):

thats all it says i think its an error. If the total number of fruit was told then i would be able to solve the question

OpenStudy (danjs):

or simply solve for # oranges in terms of apples 5x + 3y = 85 y = (85 - 5x) / 3 for a given x apples, you will have y oranges in the bag,

OpenStudy (danjs):

assume whole fruit only maybe, whole numbers x and y